# MoORE: SVD-based Model MoE-ization for Conflict- and Oblivion-Resistant Multi-Task Adaptation

### v3

@@ 4.2 Performance in Conflict-Resistance
Tables 3 and 4 compare MoORE with its competitors on CSR-MTL and NLU-MTL, respectively. Using a comparable number of learnable parameters for each dataset, MoORE achieves the best or comparable performance across all tasks and thus obtains the best average performance. These results demonstrate the superiority of MoORE in mitigating task conflict.
The impact of orthogonal adapter. In the experiment on CSR-MTL, we increase the number of Householder reflections in $\bm{H}$ (i.e., $L$) from 00 to $8$ and find that MoORE exhibits consistent improvements in performance. In the experiment on NLU-MTL, however, applying the orthogonal adapter may not improve performance. In our opinion, this phenomenon indicates that the commonsense reasoning tasks in CSR-MTL require more domain knowledge not covered by the pre-trained model. As a result, introducing the orthogonal adapter increases the number of learnable parameters and enhances the model capacity accordingly. In contrast, the text classification tasks in NLU-MTL rely more on the non-specific natural language knowledge captured by the pre-trained model. Therefore, without introducing more learnable parameters, adjusting the singular values of the pre-trained weight matrix is sufficient to achieve encouraging performance.
(a) MMLU
(b) IFEval
(c) BBH
(d) GPQA
(e) HumanEval
(f) MBPP
(g) GSM-8K
(h) Overall degradation
Figure 2: The loss of performance in OR-MTL. (a-g) The gray dashed line represents the performance of LLaMA-3.1 8B before adaptation. (h) The overall performance degradation across all tasks.
Conflict-resistance regarding task number and difficulty. To further compare and analyze the conflict-resistance capabilities of different methods, we conduct comparative experiments on CSR-MTL by varying the number and difficulty of the tasks. In particular, for each task in CSR-MTL, we first calculate the average of all the methods’ performance based on the results in Table 3. The average performance of the methods in a task measures the difficulty of the task for the base model — the lower the average performance is, the more difficult the task is. Then, we sort the tasks in ascending order based on their difficulty. Finally, we adapt the base model for the top-$K$ tasks, $K=2,...,9$, and show the performance of different adaptation methods in Figure 1(b). With the increase of task number and difficulty, all the methods suffer performance degradation because $i)$ task conflict becomes severe as the number of tasks increases, and $ii)$ difficult tasks are more likely to have conflicts with other tasks in general. Notably, MoORE consistently outperforms all other baselines across all settings.
@@ 4.3 Performance in Oblivion-Resistance
To demonstrate the superiority of MoORE in mitigating task oblivion, we compare various methods in the Language Model Evaluation Harness framework [19]. In particular, we first adapt the base model on CSR-MTL using different methods. Then, we evaluate the adapted models on OR-MTL, comparing them with the base model. Figure 2 shows that MoORE consistently mitigates task oblivion across all seven tasks of OR-MTL, with an average performance drop of only 1.31% compared to the base model. It significantly outperforms the other adaptation methods. Notably, on the HumanEval task, MoORE achieves performance exceeding the original model, with an improvement of 9.75%. Similarly, LoRAMoE and MTL-LoRA also obtain slight improvements of 4.27% and 0.16%, respectively. This intriguing phenomenon may imply that the datasets in CSR-MTL have some relevance to HumanEval, providing information and capabilities beneficial for solving this task.
Oblivion-resistance regarding task number and difficulty. We conduct experiments to investigate how the ability of oblivion-resistance changes with increased task number and difficulty. The results in the Figure 1(c) show that MoORE consistently outperforms other baselines when the number of tasks exceeds three. This result demonstrates that MoORE has stronger oblivion-resistance than its competitors. In addition, we observe an interesting phenomenon: when the number of tasks increases from 1 to 3, the performance of all methods consistently improves. However, when the number of tasks exceeds 3, their performance no longer follows a clear pattern. This may be because, when the number of tasks is no more than 3, there is insufficient diversity among the fine-tuning tasks, which leads to overfitting issues in the model.
@@
The impact of task correlation. We investigate the reason for MoORE’s oblivion-resistance empirically. In particular, we consider the samples of several sub-tasks in MMLU and those of six tasks in CSR-MTL. For each task/sub-task, we compute the average weights of MoORE’s experts over its samples, i.e., $\bm{g}_{k}=\frac{1}{|\mathcal{D}_{k}|}\sum_{\bm{x}^{(k)}\in\mathcal{D}_{k}}g(\bm{x}^{(k)})$. Given a sub-task of MMLU and a task of CSR-MTL, denoted as $\mathcal{D}_{k}$ and $\mathcal{D}_{k^{\prime}}$, respectively, we measure their correlation by $\|\bm{g}_{k}-\bm{g}_{k^{\prime}}\|_{2}$. For each sub-task of MMLU, we record the performance degradation of MoORE compared to the base model. Figure 3 shows normalized performance degradation and task correlation. This visualization result indicates that the oblivion-resistance arises from the correlation of tasks — for correlated tasks, the model can learn some common domain knowledge during adaptation and thus avoid catastrophic forgetting. In addition, this experiment also explains the result in Figure 1(c). In particular, the more tasks considered in the adaptation phase, the more likely some tasks correlate with those covered in the pre-training phase. As a result, increasing the number of tasks in the adaptation phase helps enhance the oblivion-resistance of MoORE.
@@ 5 Conclusion and Future Work
In this paper, we propose a simple but effective model MoE-ization strategy for multi-task adaptation, leading to a novel MoE architecture, called MoORE. Experiments show the effectiveness and superiority of MoORE, demonstrating that imposing orthogonality on experts and maintaining the column space of the pre-trained weight matrix help improve the resistance of the adapted model to task conflict and oblivion.
Limitations and future work. As shown in Figure 1(d), the inference efficiency of MoORE is comparable to that of baselines. However, further reducing its complexity for large-scale applications is challenging. In particular, MoORE applies many more experts than classic MoE models, and the analytic experiments in Appendix C.3 show that it learns dense weights for the experts. In addition, besides imposing orthogonal adapters, we need to find a more effective solution to enhance model capacity in broader scenarios. In theory, the results in Figures 1(b) and 1(c) show that increasing the number of tasks leads to a higher risk of task conflict while helping mitigate task oblivion. How to understand this “trade-off”? How can the limitation in multi-task adaptation be quantified regarding the number of tasks? These are still open problems. Based on the above analysis, we plan to explore efficient implementations of MoORE and study its theoretical properties in the future.
@@ Appendix C More Experimental Results
C.1 Numerical Performance in Oblivion-Resistance
Following LLaMA-3.1 8B, we select MMLU [23, 22], IFEval [76], BIG-Bench Hard (BBH) [54], GPQA [48], HumanEval [7], MBPP [4], and GSM-8K [10] as evaluation tasks and adopt the same few-shot settings. More detailed results of the oblivion-resistance experiments are presented in Table 9, corresponding to the results shown in Figures 1(c) and 2.
@@ C.2 Ablation Study
[…]
Method $D_{t}$ $D_{s}$ $L$Params Overall
[…]
@@
C.3 Analysis of Routing Weights
To investigate the router’s preferences for the experts in MoORE, we analyze the distribution of routing weights across different tasks in the CSR-MTL dataset. Take the $Q$ module from the first attention layer of the adapted model as an example. For each task, we compute and show the mean (red point) and variance (blue stem) of each expert’s weight over all samples of the task in Figures 4 and 6. In each figure, the mean weight of an expert is shown as a red point, and the blue stem associated with the point shows the standard deviation. These visualization results provide us with some insights in the mechanism behind MoORE:
@@ C.3 Analysis of Routing Weights
(a) OBQA
(b) PIQA
(c) SIQA
(d) CSQA
Figure 5: The mean and variance of routing weights obtained in four QA tasks.
(a) ARC-C
(b) ARC-E
(c) BoolQ
(d) WinoG
Figure 6: The mean and variance of routing weights obtained in the remaining four tasks of CSR-MTL.

### v4

@@ 4.2 Performance in Conflict-Resistance
Tables 5 and 6 compare MoORE with its competitors on CSR-MTL and NLU-MTL, respectively. Using a comparable number of learnable parameters for each dataset, MoORE achieves the best or comparable performance across all tasks and thus obtains the best average performance. These results demonstrate the superiority of MoORE in mitigating task conflict.
The impact of orthogonal adapter. In the experiment on CSR-MTL, we increase the number of Householder reflections in $\bm{H}$ (i.e., $L$) from 0 to $8$ and find that MoORE exhibits consistent improvements in performance. In the experiment on NLU-MTL, however, applying the orthogonal adapter may not improve performance. In our opinion, this phenomenon indicates that the commonsense reasoning tasks in CSR-MTL require more domain knowledge not covered by the pre-trained model. As a result, introducing the orthogonal adapter increases the number of learnable parameters and enhances the model capacity accordingly. In contrast, the text classification tasks in NLU-MTL rely more on the non-specific natural language knowledge captured by the pre-trained model. Therefore, without introducing more learnable parameters, adjusting the singular values of the pre-trained weight matrix is sufficient to achieve encouraging performance.
Conflict-resistance regarding task number and difficulty. To further compare and analyze the conflict-resistance capabilities of different methods, we conduct comparative experiments on CSR-MTL by varying the number and difficulty of the tasks. In particular, for each task in CSR-MTL, we first calculate the average of all the methods’ performance based on the results in Table 5. The average performance of the methods in a task measures the difficulty of the task for the base model — the lower the average performance is, the more difficult the task is. Then, we sort the tasks in ascending order based on their difficulty. Finally, we adapt the base model for the top-$K$ tasks, K=2,,9K=2,...,9, and show the performance of different methods in Figure LABEL:fig:increment1. With the increase of task number and difficulty, all the methods suffer performance degradation because $i)$ task conflict becomes severe as the number of tasks increases, and ii)ii) difficult tasks are more likely to have conflicts with other tasks. Notably, MoORE outperforms all other baselines across all settings.
(a) The loss of performance in OR-MTL. (a-g) The gray dashed line represents the performance of LLaMA-3.1 8B before adaptation. (h) The overall performance degradation across all tasks.
@@ 4.3 Performance in Oblivion-Resistance
To demonstrate the superiority of MoORE in mitigating task oblivion, we compare various methods in the Language Model Evaluation Harness framework [19]. In particular, we first adapt the base model on CSR-MTL using different methods. Then, we evaluate the adapted models on OR-MTL, comparing them with the base model. Figure 12(a) shows that MoORE consistently mitigates task oblivion across all seven tasks of OR-MTL, with an average performance drop of only 1.31% compared to the base model. It significantly outperforms the other adaptation methods. On HumanEval, MoORE achieves performance exceeding the original model, with an improvement of 9.75%. Similarly, LoRAMoE and MTL-LoRA also obtain slight improvements of 4.27% and 0.16%, respectively. This intriguing phenomenon may imply that the datasets in CSR-MTL have some relevance to HumanEval, providing information and capabilities beneficial for solving this task.
Oblivion-resistance regarding task number and difficulty. Furthermore, we conduct experiments to investigate how the ability of oblivion-resistance changes with increased task number and difficulty. The results in Figure LABEL:fig:increment2 show that after adapting to the top-3 tasks of CSR-MTL, MoORE and its strongest competitor, LoRAMoE [14], achieve comparable performance on OR-MTL. However, with the increase of task number and difficulty, the performance of LoRAMoE changes slightly while the performance of MoORE increases and becomes superior to that of LoRAMoE. This result demonstrates that MoORE has stronger oblivion-resistance than its competitors.
The impact of task correlation. We investigate the reason for MoORE’s oblivion-resistance empirically. In particular, we consider the samples of several sub-tasks in MMLU and those of six tasks in CSR-MTL. For each task/sub-task, we compute the average weights of MoORE’s experts over its samples, i.e., $\bm{g}_{k}=\frac{1}{|\mathcal{D}_{k}|}\sum_{\bm{x}^{(k)}\in\mathcal{D}_{k}}g(\bm{x}^{(k)})$. Given a sub-task of MMLU and a task of CSR-MTL, denoted as $\mathcal{D}_{k}$ and $\mathcal{D}_{k^{\prime}}$, respectively, we measure their correlation by $\|\bm{g}_{k}-\bm{g}_{k^{\prime}}\|_{2}$. For each sub-task of MMLU, we record the performance degradation of MoORE compared to the base model. Figure 12 shows normalized performance degradation and task correlation. This visualization result indicates that the oblivion-resistance arises from the correlation of tasks — for correlated tasks, the model can learn some common domain knowledge during adaptation and thus avoid catastrophic forgetting.
The generalization to unseen task IDs. The computation of task-level weights requires task IDs as input, but the ID of a new task is unavailable during inference. Therefore, we use the average of the existing task ID embeddings as the embedding of the new task. Specifically, after adapting the base model on CSR-MTL, MoORE only learns the task IDs of the nine tasks in CSR-MTL. During evaluation on OR-MTL, we use the average of the nine task ID embeddings as the task ID embedding of each new task. The results in Figure LABEL:fig:increment2 and 12(a) validate the effectiveness of this approach.
In addition, this experiment also explains the result in Figure LABEL:fig:increment2. In particular, the more tasks considered in the adaptation phase, the more likely some tasks correlate with those covered in the pre-training phase. As a result, increasing the number of tasks in the adaptation phase helps enhance the oblivion-resistance of MoORE.
4.4 Inference Efficiency
To compare inference efficiency, we adapt the base model using MoORE and other baselines, and evaluate their inference time on GPQA. For a fair comparison, all experiments are conducted on a single NVIDIA A100 GPU with the same batch size. The results in Figure LABEL:fig:speed show the average inference time per batch for each method, which indicates that MoORE achieves superior inference efficiency compared to most baselines and is comparable to MixLoRA. Compared to the base model (i.e., the gray dashed line), MoORE moderately increases inference time.
@@ 5 Conclusion and Future Work
In this paper, we propose a simple but effective model MoE-ization strategy for multi-task adaptation, leading to a novel MoE architecture, called MoORE. Experiments on various datasets show the effectiveness and superiority of MoORE, demonstrating that imposing orthogonality on experts and maintaining the column space of the pre-trained weight matrix help improve the resistance of the adapted model to task conflict and oblivion.
Limitations and future work. As shown in Figure LABEL:fig:speed, the inference efficiency of MoORE is comparable to that of baselines. However, further reducing its complexity for large-scale applications is challenging. In particular, MoORE applies many more experts than classic MoE models, and the analytic experiments in Appendix C.4 show that it learns dense weights for the experts. In addition, besides imposing orthogonal adapters, we need to find a more effective solution to enhance model capacity in broader scenarios. In theory, the results in Figures LABEL:fig:increment1 and LABEL:fig:increment2 show that increasing the number of tasks leads to a higher risk of task conflict while helping mitigate task oblivion. How to understand this “trade-off”? How can the limitation in multi-task adaptation be quantified regarding the number of tasks? These are still open problems. Based on the above analysis, we plan to explore efficient implementations of MoORE and study its theoretical properties in the future.
Acknowledgments and Disclosure of Funding
This work was supported by National Natural Science Foundation (92270110), Beijing Natural Science Foundation (L233008), the Fundamental Research Funds for the Central Universities, and the Research Funds of Renmin University of China. We also acknowledge the support provided by the fund for building world-class universities (disciplines) of Renmin University of China and by the funds from Engineering Research Center of Next-Generation Intelligent Search and Recommendation, Ministry of Education, and from Intelligent Social Governance Interdisciplinary Platform, Major Innovation & Planning Interdisciplinary Platform for the “Double-First Class” Initiative, Renmin University of China.
@@ Appendix C More Experimental Results
C.1 Performance on other Model Architectures in Conflict-Resistance
To evaluate the cross-model performance of MoORE, we take Qwen-2.5 7B as the backbone model and apply various MoE-based multi-task adaptation methods to it. Experimental results on CSR-MTL are shown in Table 11. We can find that MoORE outperforms baselines consistently under different hyperparameter settings, and the performance of MoORE is robust to the change of $L$. These results verify the robustness of our method and its compatibility with various model architectures.
Table 11: Results (%) of Qwen-2.5 7B adapted by various methods on CSR-MTL. The best results on each dataset are shown in bold, and the second best results are shown in underline.
LoRA [25] 2.12% 87.29 91.88 70.37 90.60 87.92 79.02 92.73 74.98 84.36 84.35
MixLoRA [30] 3.32% 86.86 91.75 71.38 91.40 89.28 80.45 95.06 78.45 85.09 85.53
MoSLD [81] 1.43% 86.69 92.13 71.96 90.40 89.66 80.40 95.17 78.30 85.50 85.58
HydraLoRA [60] 2.83% 88.23 91.54 72.57 91.20 85.80 80.09 95.95 82.24 85.26 85.87
MTL-LoRA [73] 2.80% 87.80 92.09 72.72 92.20 86.07 79.79 95.92 80.66 84.36 85.73
LoRAMoE [14] 2.21% 87.20 91.84 73.61 90.40 90.37 81.27 95.99 80.27 84.60 86.17
[…]
C.2 Numerical Performance in Oblivion-Resistance
Following LLaMA-3.1 8B, we select MMLU [24, 23], IFEval [83], BIG-Bench Hard (BBH) [58], GPQA [51], HumanEval [7], MBPP [4], and GSM-8K [10] as evaluation tasks and adopt the same few-shot settings. More detailed results of the oblivion-resistance experiments are presented in Table 12, corresponding to the results shown in Figures LABEL:fig:increment2 and 12(a).
@@ C.3 Ablation Study
Table 13: Ablation studies of MoORE on CSR-MTL.   indicates the absence of that component.
Method $D_{t}$ $D_{s}$ $L$ #Params Overall
16 0.14% 79.55
Merely Apply 32 0.29% 81.80
Task-level Routing 64 0.58% 83.06
128 1.16% 83.53
16 0.39% 82.30
Merely Apply 32 0.78% 82.91
Sample-level Routing 64 1.57% 83.57
128 3.13% 83.80
2 0.03% 80.76
Merely Apply 4 0.06% 82.15
Orthogonal Adapter 8 0.12% 82.98
16 0.25% 83.90
@@
C.4 Analysis of Routing Weights
To investigate the router’s preferences for the experts in MoORE, we analyze the distribution of routing weights across different tasks in the CSR-MTL dataset. Take the $Q$ module from the first attention layer of the adapted model as an example. For each task, we compute and show the mean (red point) and variance (blue stem) of each expert’s weight over all samples of the task in Figures 13 and 17(b). In each figure, the mean weight of an expert is shown as a red point, and the blue stem associated with the point shows the standard deviation. These visualization results provide us with some insights in the mechanism behind MoORE:
@@ C.4 Analysis of Routing Weights
(a) The mean and variance of routing weights obtained in four QA tasks.
(b) The mean and variance of routing weights obtained in the remaining four tasks of CSR-MTL.